Migrants are moving from a mainland population to an island at an average rate of 50 individuals per generation. The frequency of A2 is 80% on the mainland while 0% of the 1000 individuals currently on the island carry allele A2. Allele A2 turns out to have a selective advantage, such that homozygous individuals reproduce at four times the rate of the native population What is the frequency q of A2 individuals on the island after selection, but before migration What is the frequency q of A2 after selection and migration? What is the frequency q of A2 individuals on the island after one more round of selection, but before any more migration? For all answers (please use this format with 3 significant figures e.g. for an answer of 0.5: 0.500)
Answer #1

According to Hardy-Weinberg Equilibrium:

   ?+?= 1

The frequency q* of individuals on the island after selection but before migration = 0.000

Because 0% of individuals on island carry A2 allele.

Frequency q of A2 after selection and migration = number of migrants/ total population

                                                                                        = 50/1050 = 0.047

The frequency q* of individuals on the island after one more round of selection, but before anymore migration = q * 4 = 0.047 * 4 = 0.188 .

Because Allele A2 has a selective advantage of reproducing 4 times faster than native population